OpenStudy (anonymous):

simplify 10^3(10^5+10^5) / 10^4

jhonyy9 (jhonyy9):

\[2*10^{5}/10 =2*10^{4}\]

jhonyy9 (jhonyy9):

because

jhonyy9 (jhonyy9):

\[10^{3}/10^{4} = 1/10\]